# RooflineBench: A Benchmarking Framework for On-Device LLMs via Roofline Analysis

### v1

@@ 7 Conclusion
We characterize LLM inference on edge devices using the Roofline model and introduce Relative Inference Potential to compare efficiency across heterogeneous substrates. Findings show that sequence length dictates distinct performance and operational intensity ($OI$) dynamics. We identify a non-monotonic $OI$ trajectory: performance initially rises via overhead amortization but regresses beyond 3–5 layers as parameter loading outpaces computational reuse. Our analysis reveals an efficiency trap from hardware ridge-point disparities, while structural refinements sustain higher $OI$ via optimized cache management. These results advocate for hardware-software co-design to align model architectures with physical constraints and unlock localized intelligence.
@@ Appendix B Future Work
While this study provides an extensive empirical characterization of mainstream small language models across diverse hardware tiers, we identify several strategic avenues for future exploration. Primarily, we intend to expand the taxonomy of our benchmarking framework to incorporate a broader range of attention mechanisms. Although our current analysis provides a comprehensive evaluation of critical patterns such as $MLA$ and $GQA$, we aim to adapt our methodology to emerging structural variants and hybrid architectures to ensure the continued relevance of our operational intensity ($OI$) insights across the evolving architectural landscape.
A significant frontier lies in the analytical characterization of Mixture of Experts (MoE) architectures. Quantifying the theoretical $FLOPS$ for MoE models during dynamic inference remains a formidable challenge because the stochastic nature of token routing complicates the estimation of the execution ceiling. While the current work focuses on dense architectures, extending the Roofline model to account for sparse activation regimes will be essential for identifying the unique physical constraints and efficiency traps inherent in localized MoE deployment.
Finally, we plan to broaden the experimental scope across both hardware and software dimensions. On the hardware side, we aim to scale our testing to a more diverse array of heterogeneous edge devices to further validate the cross platform robustness of our findings. On the software side, we recognize that different inference engines, such as TensorRT LLM, vLLM, and ONNX Runtime, introduce substantial variance in performance. Investigating how these software stacks influence $OI$ and realized throughput will provide a more holistic understanding of the entire deployment pipeline, facilitating a more effective transition toward true hardware software co design.
@@ C.3 Experimental Setting
FLOPs Calculation Methodology. To evaluate operational intensity ($OI$) and performance within the Roofline model, we estimate the floating point operations ($FLOPs$) associated with each inference configuration. This estimation framework accounts for linear projections, attention mechanisms, and KV cache management by leveraging key architectural parameters including the hidden dimension ($h$), the number of query heads ($n_{h}$), and the number of key-value heads ($n_{kv}$). The specific estimation logic Sun et al. (2025), with detailed formulas provided in Table 2. The resulting $FLOPs$ values serve as an approximated baseline for deriving $OI$ and throughput metrics.
@@ Appendix D Comprehensive Benchmarking Analysis of Model Architectures on Heterogeneous Hardware Platforms
This appendix provides a comprehensive suite of benchmarking results to substantiate the empirical findings and theoretical analyses presented in the main text. We detail the performance profiles of various large language model architectures including Multi head Latent Attention and trainable sparse attention across a spectrum of heterogeneous hardware platforms ranging from high performance accelerators to edge side SoCs. The presented data encompass a wide array of quantization precisions and operational scenarios while offering granular insights into the relationship between architectural design, layer wise operational intensity ($OI$), and hardware resource saturation. By providing these extended datasets, we aim to offer a rigorous validation of the scaling dynamics and utilization equity discussed throughout the study.

### v2

@@ 7 Conclusion
We characterize LLM inference on edge devices using the Roofline model and introduce Relative Inference Potential to compare efficiency across heterogeneous substrates. Findings show that sequence length dictates distinct performance and operational intensity (OIOI) dynamics. We identify a non-monotonic OIOI trajectory: performance initially rises via overhead amortization but regresses beyond 3–5 layers as parameter loading outpaces computational reuse. Our analysis reveals an efficiency trap from hardware ridge-point disparities, while structural refinements sustain higher OIOI via optimized cache management. These results advocate for hardware-software co-design to align model architectures with physical constraints and unlock localized intelligence.
@@ Appendix B Future Work
While this study provides an extensive empirical characterization of mainstream small language models across diverse hardware tiers, we identify several strategic avenues for future exploration. Primarily, we intend to expand the taxonomy of our benchmarking framework to incorporate a broader range of attention mechanisms. Although our current analysis provides a comprehensive evaluation of critical patterns such as MLAMLA and GQAGQA, we aim to adapt our methodology to emerging structural variants and hybrid architectures to ensure the continued relevance of our operational intensity (OIOI) insights across the evolving architectural landscape.
A significant frontier lies in the analytical characterization of Mixture of Experts (MoE) architectures. Quantifying the theoretical FLOPSFLOPS for MoE models during dynamic inference remains a formidable challenge because the stochastic nature of token routing complicates the estimation of the execution ceiling. While the current work focuses on dense architectures, extending the Roofline model to account for sparse activation regimes will be essential for identifying the unique physical constraints and efficiency traps inherent in localized MoE deployment.
Finally, we plan to broaden the experimental scope across both hardware and software dimensions. On the hardware side, we aim to scale our testing to a more diverse array of heterogeneous edge devices to further validate the cross platform robustness of our findings. On the software side, we recognize that different inference engines, such as TensorRT LLM, vLLM, and ONNX Runtime, introduce substantial variance in performance. Investigating how these software stacks influence OIOI and realized throughput will provide a more holistic understanding of the entire deployment pipeline, facilitating a more effective transition toward true hardware software co design.
@@ C.3 Experimental Setting
FLOPs Calculation Methodology. To evaluate operational intensity (OIOI) and performance within the Roofline model, we estimate the floating point operations (FLOPsFLOPs) associated with each inference configuration. This estimation framework accounts for linear projections, attention mechanisms, and KV cache management by leveraging key architectural parameters including the hidden dimension ($h$), the number of query heads ($n_{h}$), and the number of key-value heads ($n_{kv}$). The specific estimation logic Sun et al. (2025), with detailed formulas provided in Table 2. The resulting FLOPsFLOPs values serve as an approximated baseline for deriving OIOI and throughput metrics.
@@ Appendix D Comprehensive Benchmarking Analysis of Model Architectures on Heterogeneous Hardware Platforms
This appendix provides a comprehensive suite of benchmarking results to substantiate the empirical findings and theoretical analyses presented in the main text. We detail the performance profiles of various large language model architectures including Multi head Latent Attention and trainable sparse attention across a spectrum of heterogeneous hardware platforms ranging from high performance accelerators to edge side SoCs. The presented data encompass a wide array of quantization precisions and operational scenarios while offering granular insights into the relationship between architectural design, layer wise operational intensity (OIOI), and hardware resource saturation. By providing these extended datasets, we aim to offer a rigorous validation of the scaling dynamics and utilization equity discussed throughout the study.